# Redundancy of the cosmological evolution equations and its relationship with the initial conditions

### v2

@@ 5 On the logical consistency of the theory
As the Einstein equations (Friedmann and pressure equations) in cosmology do not contain a time derivative of the energy density of the fluid so those equations does not generate naturally a differential equation for $\rho(t)$, but we know from general assumptions, about the system, that such an equation is inevitable. In such a case there is only one way the first order differential equation for $\rho(t)$ can appear, that intended equation must be derivable from the available Einstein equations. In reality it happens like that as because the continuity equation is obtained from the Friedmann and pressure equations. To obtain the continuity equation one must have to differentiate the Friedmann equation and consequently the Friedmann equation must be a first order differential equation. The Friedmann equation cannot be a second order differential equation as in that case a differentiation of this equation will produce a third order derivative of the scale-factor in the resulting equation which cannot be cancelled by any terms in the other Einstein equation for pressure. Hence the equation giving the energy density of the fluid in terms of the functions of the scale-factor and its derivatives must have to be a first-order differential equation of $a(t)$. As because GR, as a well defined field theory, contains second order derivative terms of the scale-factor, the other equation for pressure must have to be a second order differential equation for $a(t)$.
At this point we must point out that with all the manipulations, as specified in the previous paragraphs, one in general always obtains the desired first order differential equation for $\rho(t)$. This is not a coincidence, rather the theory is formulated in such a way that the desired differential equation is produced due to a tensor identity, called the (second) Bianchi identity, constraining the Einstein tensor. The Bianchi identity naturally makes the theory consistent by introducing the energy conservation condition as predicted by the continuity equation.
The previous discussion made it clear that in GR, the Bianchi identity and the Einstein equation both are responsible for the dynamical structure of the cosmological theory and together they affect the initial condition of the problem. Interestingly, whatever we have discussed about cosmological dynamics in GR can be generalized to a modern modified theory of gravity. In the next section we generalize our results in $f(R)$ gravity.
6 On the redundancy of dynamical equations and initial conditions in $f(R)$ cosmology
Till now we have worked in the ambit of GR and Newtonian cosmology. In this section we will show that the techniques we have derived can be generalized in a modified theory of gravity, called the metric $f(R)$ theory. This modification of gravity is a modern development which became useful in theories of inflation and tackling the dark energy problem. The two reviews in Refs. [14, 15] will help the reader to grasp the essence of the theory and we will closely follow the conventions and notations used in those reviews.
In metric $f(R)$ gravity theory the gravitational action changes from
$S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,R+S_{m}\to S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,f(R)+S_{m}\,$,
where $R$ is the Ricci scalar and $S_{m}$ is the action for the non-gravitational degrees of freedom, generally called the matter action. Here $f(R)$ is a general function of $R$ and we do not choose any particular form of $f(R)$ here as our analysis will be true for any general $f(R)$. The field equations corresponding to the Eqs. (3), (4) in $f(R)$ gravity turns out to be:
$\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $\frac{8\pi G}{3f^{\prime}(R)}(\rho+\rho_{\rm{eff}})\,$, (24)
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $-\frac{8\pi G}{f^{\prime}(R)}(P+P_{\rm{eff}})\,$, (25)
where a prime denotes a derivative with respect to the Ricci scalar, $R$, and
$8\pi G\rho_{\rm{eff}}$ $=$ $\frac{Rf^{\prime}(R)-f(R)}{2}-3H\dot{R}f^{\prime\prime}(R)\equiv\tilde{\rho}_{\rm{eff}}\,$, (26)
$8\pi GP_{\rm{eff}}$ $=$ $\dot{R}^{2}f^{\prime\prime\prime}(R)+2H\dot{R}f^{\prime\prime}(R)+\ddot{R}f^{\prime\prime}(R)-\frac{Rf^{\prime}(R)-f(R)}{2}\equiv\tilde{P}_{\rm{eff}}\,$, (27)
where we have used the Hubble parameter, $H\equiv\dot{a}/a$, to compactly write the above two expressions. The Ricci scalar $R=6[(\ddot{a}/{a})+(\dot{a}^{2}/{a^{2}})+({k}/{a^{2}})]=6[\dot{H}+2H^{2}+(k/a^{2})]$ for the FLRW metric given in Eq. (1). Here $\rho_{\rm{eff}},P_{\rm{eff}}$ are effectively acting like the energy density and the pressure induced by the Ricci curvature of the spacetime. In the above equations $\rho,P$ are the normal energy density and pressure of a barotropic matter fluid present in the FLRW spacetime. We name Eq. (24) as the modified Friedmann equation and Eq. (25) as the modified pressure equation. One can also combine the field equations to write down the modified acceleration equation in the present case. Although we name the relevant field equations as we did before, one must remember that these equations are very different from the corresponding equations in GR as here the modified Friedmann equation is a third order differential equation (apparent from the expression for $\rho_{\rm eff}$), the modified pressure equation and the modified acceleration equation are fourth order differential equations (apparent from the expression for $P_{\rm eff}$).
Remarkably, in the present case if one differentiates the modified Friedmann equation and combines the result with the pressure equation, one gets back the continuity equation for the barotropic fluid, as given in Eq. (6). Consequently, in the present case we have four differential equations of three different orders: one is a third order equation, two are fourth order equations and the other one is a first order equation. The disparity in the degrees of the dynamical equations has increased in the present case. Interestingly, with all these drastic differences with GR, we will show that most of the techniques used to understand the redundancy of the equations are still applicable in the present case and the special role of the modified Friedmann equation remains intact in $f(R)$ gravity.
The above equations show that $f^{\prime}(R)$ modifies the gravitational coupling terms from $G\to G/f^{\prime}(R)$, consequently for gravitational stability of the theory we demand $f^{\prime}(R)>0$. Moreover, it was noted that if $f^{\prime\prime}(R)$ is not positive, one can have another kind of instability of the theory, called the Dolgov-Kawasaki instability [16]. As a consequence of these facts, we will always assume
$f^{\prime}(R)>0\,,\,\,\,\,\,\,\,\,f^{\prime\prime}(R)>0\,$, (28)
in $f(R)$ cosmology. In the field equations of $f(R)$ gravity we have set the cosmological constant, $\Lambda=0$, as $f(R)$ theory can dynamically generate dark energy and we may not explicitly require a cosmological constant in this framework of gravity.
In the present case one can differentiate the modified Friedmann equation with respect to time and combine the result with the continuity equation to obtain the pressure equation. We do not present this derivation as it is a straightforward calculation. We show how one obtains the modified Friedmann constraint from the modified pressure equation and the continuity equation. This calculation will make it clear how the previous concepts can be generalized in the present case. If we choose Eq. (25) and the continuity equation as our basic dynamical equations then we can start to manipulate them in a way done previously. Using the continuity equation we write Eq. (25) as:
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}+\frac{8\pi G}{f^{\prime}}P_{\rm eff}=\frac{8\pi G}{f^{\prime}}\left(\frac{1}{3}\dot{\rho}\frac{a}{\dot{a}}+\rho\right)\,.$ (29)
Multiplying all the terms of the above equation by $\dot{a}/a$ we get
$2\frac{\ddot{a}\dot{a}}{a^{2}}+\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\dot{\rho}+\frac{8\pi G}{f^{\prime}}\rho\frac{\dot{a}}{a}\,.$ (30)
The above equation can also be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}\right)+\frac{3\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\dot{\rho}+\frac{8\pi G}{f^{\prime}}\rho\frac{\dot{a}}{a}\,.$
Subtracting
$3\left(\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}\right)$
from both sides of the last equation, we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\frac{d\rho}{dt}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3f^{\prime}}\rho-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,$, (31)
which can also be written as:
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}$ $=$ $\frac{d}{dt}\left[\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})\right]-\frac{8\pi G}{3f^{\prime}}\frac{d\rho_{\rm eff}}{dt}+\frac{8\pi G}{3f^{\prime\,2}}(\rho+\rho_{\rm eff})f^{\prime\prime}\dot{R}$ (32)
- $\frac{8\pi G}{f^{\prime}}H\rho_{\rm eff}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,.$
Defining
$Q(t)\equiv\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)-\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})\,$, (33)
we can write Eq. (32) as:
$\frac{dQ}{dt}+3HQ=-\frac{1}{3f^{\prime}}\left[\frac{d\tilde{\rho}_{\rm eff}}{dt}+3H(\tilde{\rho}_{\rm eff}+\tilde{P}_{\rm eff})-3\left\{\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})f^{\prime\prime}\dot{R}\right\}\right]\,.$ (34)
In the present case one can show that:
$\frac{d\tilde{\rho}_{\rm eff}}{dt}+3H(\tilde{\rho}_{\rm eff}+\tilde{P}_{\rm eff})=\dot{R}f^{\prime\prime}\left(\frac{1}{2}R-3\dot{H}-3H^{2}\right)=3\dot{R}f^{\prime\prime}\left(H^{2}+\frac{k}{a^{2}}\right)\,$, (35)
which show that the energy density for the curvature term does not follow a conservation equation. Moreover the above equation reveals that only when $R$ is a constant one may expect that the effective curvature energy density and pressure will satisfy the hydrodynamic equation of continuity. $R$ can turn out to be a constant when $k=0$ and $H$ is a constant, specifying a de Sitter universe. Using the above result we can write Eq. (34) as:
$\frac{dQ}{dt}+3\left(H+\frac{f^{\prime\prime}\dot{R}}{3f^{\prime}}\right)Q=0\,$, (36)
which shows that if $Q=0$ initially, then $Q$ will always be zero. This shows that if we choose the initial conditions to be satisfying the new Friedmann equation then, while working with the new pressure equation and the continuity equation, the new Friedmann equation will always be satisfied.
The above calculation shows that in the present theory, which is very different from GR, we can ask the same questions we asked before. The questions are related to the redundancy of dynamical equations. The redundancy problem ultimately leads to the special status of the modified Friedmann equation. This workout shows that in $f(R)$ cosmology all the initial conditions must have to satisfy the modified Friedmann equation which acts like a constraint equation.
Fortunately, in the present case there exists a particular conformal transformation, which converts the whole problem into a form where the gravitational dynamics is generated by the Einstein equations in GR. Once we convert the problem into a problem of GR, we can then apply our understanding of the initial value problem in GR to unravel the initial value problem in $f(R)$ theories. As because conformal transformations produces GR like dynamics where Einstein equations dictate gravitational dynamics, our old discussions hold in the conformally transformed frame, called the Einstein frame. In this language the system (or the frame) where the dynamics is governed by the original $f(R)$ equations (as above, in particular the Eqs. (24), (25)) is called the Jordan frame. It is seen that our general discussion about the redundancy of cosmological equations and the special role of the Friedmann equation holds true both in the Jordan frame and in the Einstein frame.
@@ 7 Conclusion
In this paper we have analyzed the dynamical structure of cosmological evolution in GR. In particular we addressed the issues related to the redundancy of cosmological evolution equations and the relationship of this redundancy with the initial value problem in cosmology. It was then shown that in a modern modified theory of gravity, called $f(R)$ gravity, a similar issue of redundancy of cosmological equations persist. Interestingly, our solution to the redundancy problem in GR sheds important light in solving the redundancy problem in $f(R)$ cosmology. There is a branch of cosmology, called Newtonian cosmology, where one obtains the basic cosmological equations in GR in a dust dominated universe primarily using some adhoc assumptions on the nature of spacetime and semi-Newtonian conceptions [17, 18, 19]. As the basic equations of cosmology in these models remain the same as in GR, one faces the same redundancy problem in these cases and our solution of the redundancy problem works in Newtonian cosmology. It is seen that by focusing our attention on the solvability of the dynamical system we can learn important properties of the Einstein equation in cosmology and unravel the special role of the Bianchi identity. Throughout we have applied a bottom-up approach and it seems that this approach yields interesting results. Our approach will remain relevant, primarily, for two different points. The first point is related to the simplicity and accessibility of the analysis. The present article shows that we can attain some interesting insights about the structure of the cosmological equations starting from very rudimentary assumptions. The second point is related to the fact that the results about the redundancy problem in cosmology in GR can be generalized to more modern variants of theory of gravity, as $f(R)$ gravity.

### v3

@@ 5 On the logical consistency of the theory
As the Einstein equations (Friedmann and pressure equations) in cosmology do not contain a time derivative of the energy density of the fluid so those equations does not naturally generate a differential equation for $\rho(t)$, but we know from general assumptions that such an equation (as predicted by the Bianchi identity) is inevitable. In such a case there is only one way the first order differential equation for $\rho(t)$ can appear, that intended equation must be derivable from the available Einstein equations. This implies that one of the Einstein equations, containing $\rho$ on the right hand side, must have to be a first order differential equation in the scale-factor, $a(t)$. It turns out that equation is the Friedmann equation. The Friedmann equation cannot be a second order differential equation, for $a(t)$, as in that case a differentiation of this equation will produce a third order derivative of the scale-factor in the resulting equation which cannot be cancelled by any terms in the other Einstein equation for pressure. As because GR, as a well defined field theory, contains second order derivative terms of the scale-factor, the other equation for pressure must have to be a second order differential equation for $a(t)$.
These arguments show the subtle relation between the redundancy problem and the special role of the Bianchi identity in GR. The presence of the Bianchi identity compels the FLRW spacetime to have different orders of differential equations which leads to the Friedmann constraint for fixing equivalent initial conditions.
@@ 6 Conclusion
In this paper we have analyzed the dynamical structure of cosmological evolution in GR. In particular we addressed the issues related to the redundancy of cosmological evolution equations and the relationship of this redundancy with the initial value problem in cosmology. There is a branch of cosmology, called Newtonian cosmology, where one obtains the basic cosmological equations in GR in a dust dominated universe primarily using some ad-hoc assumptions on the nature of spacetime and semi-Newtonian conceptions [14, 15, 16]. As the basic equations of cosmology in these models remain the same as in GR, one faces the same redundancy problem in these cases and our solution of the redundancy problem works in Newtonian cosmology.
It is seen that by focusing our attention on the solvability of the dynamical system we can learn important properties of the Einstein equation in cosmology and unravel the special role of the Bianchi identity. Throughout we have applied a bottom-up approach and it seems that this approach yields interesting results. The strength of our approach is related to the simplicity and accessibility of the analysis.